OpenStudy (anonymous):

how do you solve this rational equation 1/x=x/9?

OpenStudy (anonymous):

\[\frac{1}{x}=\frac{x}{9}\] \[9=x^{2}\]\[x=\sqrt{9}\]\[x=3\]

OpenStudy (anonymous):

Check by back substitution: \[\frac{1}{(3)}=\frac{(3)}{9}\]\[\frac{1}{3}=\frac{1}{3}\] CHECK

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

how do you solve the rational equation 3x/4=5x+1/3